OpenStudy (anonymous):

this is my last day to do my work if i don't then i will fail math can some one please help me ?

OpenStudy (anonymous):

I hate to be Mr. obvious, but this is the wrong day to start.

OpenStudy (anonymous):

^^

OpenStudy (mathmale):

so, betty lee, please post the question you want help with, right off. Next time, don't spill the beans about having waited until the last minute; then we can't tease you.

OpenStudy (mathmale):

Thanks for your wisdom, Mr. Obvious.

OpenStudy (anonymous):

okay @mathmale thank you so much im not trying to do a sob story but ive really being crying i need to pass math but i dont get anything.

OpenStudy (anonymous):

The table shows the number of miles driven over time. Express the relationship between distance and time in simplified form as a unit rate: ___________ mph. Write a statement that interprets the relationship: _________________________________________________. Time (hours) 3 5 7 9 Distance (miles) 171 285 399 513

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Mathmale I'm stepping out, less confusion 1 on 1. betty, he will steer you straight.

OpenStudy (mathmale):

OK, Betty Lee. I won't rub it in. Next time you have a test coming up, get onto OpenStudy one full week in advance. Now, Betty, I'm going to Bet that this is a linear relationship (such as y=mx+b).

OpenStudy (mathmale):

After 3 hours the distance is 171. After 5 hours the distance is 285. What was the change in time? What was the change in distance?

OpenStudy (anonymous):

well i have been on open study for the pass two weeks no one has helped me and i don't understand and i have an f if i don't pass then i don't know what to do!

OpenStudy (mathmale):

@ehuman : All the best to you. A worthy fellow you are. I'm assuming you're a fellow, of course.

OpenStudy (anonymous):

i dont know anythign @mathmale

OpenStudy (mathmale):

Please stop saying that. Think, "with time on task and proper guidance, I can learn just about anything my teachers throw at me."

OpenStudy (anonymous):

lol yes, my avatar is a pic of me and my youngest son when he was just born.

OpenStudy (mathmale):

How much does t change? Simply subtract 3 from 5, and you end up with 2 (hours). How much does the distance change? Simply subtract 171 from 285.

OpenStudy (anonymous):

well everyone is different this is my 3rd time failing math so yeah!!!!!!

OpenStudy (mathmale):

Please, betty lee, stop the negativism; it simply hurts you.

OpenStudy (mathmale):

Please answer my question.

OpenStudy (anonymous):

114?

OpenStudy (mathmale):

Yes. Great. that's in miles. If you divide 114 miles by 2 hours, what do you get? Be certain to include the units of measurement (miles, hours).

OpenStudy (anonymous):

57 miles right

OpenStudy (mathmale):

Yes, that's fine. Very nice. Now go back to the first question I asked you: What's the change from 171 miles to 285 miles? (subtract 171 from 285).

OpenStudy (anonymous):

sorry @mathmale my computer went dead

OpenStudy (mathmale):

So did mine. So I went to take a nap! What's the change, from 171 to 285 miles?

OpenStudy (anonymous):

okay so the first question I don't understand

OpenStudy (mathmale):

You've already answered one of my questions properly. You divided 114 by 2 and got 57. If you were to do the same type of calculation with 285-171 you'd get 114 again. You divide that by 2 (hours) and you end up with 57 miles per hour, or 57 mi/hr..

OpenStudy (mathmale):

so you have essentially solved this first problem.

OpenStudy (anonymous):

okay so the answer was 57mils which I had already put down as my answer

OpenStudy (mathmale):

Go back and look at the problem that you posted. It asks you to produce the unit rate: 57 miles per hour. You already have that! Here's what it means in words: Each hour, the driver covers 57 miles. So the rate of change of distance with time is 57 miles per hour." then you're done.

OpenStudy (mathmale):

that's it. Post the next question, please.

OpenStudy (anonymous):

Find the slope of the line. M= __________

OpenStudy (anonymous):

What is the slope of the line? M= _____________

OpenStudy (mathmale):

This focuses on the same principle: rate of change. Earlier, you subtrancted mileage and subtracted time; you ended up finding the rate of change, or slope, which was 57 miles/hr. This graph gives you data in a different manner: it's graphical data. Look at those arrowheads. One of them is at (4,7) and the other at (-4,-5). You need to calculate how much y changes as you travel left to right from -5 to +5. Do it, please. You could count the number of squares between -5 and +7, or you could subtract (-5) from (+7).

OpenStudy (anonymous):

-2 ? that's what im subtracting right

OpenStudy (mathmale):

Just do this: 7-(-5)= ???? and we'll answer your verbal question later. In the third problem, y does not change. In other words, your slope is zero. Simple as that.

OpenStudy (mathmale):

what is (-)(-) ?

OpenStudy (anonymous):

12

OpenStudy (mathmale):

Important: the slope formula is m=rise/run. \[slope=m=\frac{ rise }{ run }\]

OpenStudy (mathmale):

This is what you're calculating right now for the 2nd problem. Note that this formula seems to be mentioned in both Problem 2 and Problem 3. You'll need to find some way to learn and memorize this formula.

OpenStudy (mathmale):

Yes, thank you. 7-(-5)=12. You're doing fine. Now find 5-(-5).

OpenStudy (anonymous):

okay so was that answer correct >

OpenStudy (mathmale):

Yes, it was correct. The "rise" (meaning the change in y) was 12.

OpenStudy (anonymous):

10

OpenStudy (mathmale):

Now you're calculating the "run," the change in x. Yes, it's 10. Now divide 12 by 10. that's rise/run. Simplify. What do you get?

OpenStudy (anonymous):

I got 1.2

OpenStudy (mathmale):

\[\frac{ rise }{ run}=\frac{ 12 }{ 10 }=what.fraction?\]

OpenStudy (mathmale):

Cool. That's the answer. In #2, the slope is 12/10, or 1.2. See? You're doing fine. Now in the third problem, y does not change. Thus, the rise is zero, but the run is not. What is the slope? rise/run = ???

OpenStudy (anonymous):

is it 1 over 1 ?

OpenStudy (mathmale):

Please go back and re-read what I typed. The rise is ... ? Then the slope is rise/run = ( ??? )/run = ???

OpenStudy (anonymous):

I just did I don't understand how you got the number

OpenStudy (anonymous):

is it 1/2

OpenStudy (mathmale):

rise = "new" y - "old" y. Here, y starts at -2 and ends at -2 (no change). What is -2-(-2)? rise = ????

OpenStudy (mathmale):

Please calculate the rise, as defined above.

OpenStudy (anonymous):

0

OpenStudy (mathmale):

That's right. supposing that "run" is not zero, what would slope=rise/run be equal to?

OpenStudy (anonymous):

2/2 or 2.0

OpenStudy (anonymous):

or 1

OpenStudy (mathmale):

Don't see how you got that . We want the ratio rise/run. If rise is 0 and run is not 0, what is their quotient?

OpenStudy (anonymous):

honestly I don't know if its 0 then 0 right ???????

OpenStudy (mathmale):

Yes. Why? 0/(5)=0, right? Doesn't matter what the denominator is, so long as it's not zero. 0 div by a non-zero number is 0. So, your slope is zero (0). You should be able to recognize that instantly , visually, from the graph: that horiz. line has a slope of zero.

OpenStudy (anonymous):

oh ojay that was my first answer but I didn't think it could only be 0

OpenStudy (mathmale):

It's extremely important that when you see a horiz. line, you think immediately": the slope is zero. Also, you should know that 0 divided by any nonzero quantity is 0. I ahve time for one more question. Have you more qeustions?

OpenStudy (anonymous):

What is the slope of a line passing through the pair of points (2, 10) and (8, 3) that's the next question

OpenStudy (mathmale):

You'll need to find the slope of that line connecting the two given points. You do exactly the same thing as before: you calculate m = rise / run. This problem will be a little harder because in this case the slope is negative; that means that if x increases, y decreases. Pls review what we did earlier and see if you can find rise/run. (there's more than one way to do it!!)

OpenStudy (anonymous):

8/-5

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@nikato can you help

OpenStudy (nikato):

No, that's not correct. Try using the slope formula

OpenStudy (nikato):

http://www.purplemath.com/modules/slope.htm

OpenStudy (nikato):

So if u substitute in ur points to find the slope, u'll get 10-3 ------= 2-8 Can u simplify this?

OpenStudy (mathmale):

Betty: as I suggested before, the result of these calculations will be negative; you have a negative slope, a negative "rise" with a positive "run" Memorize that slope formula.

OpenStudy (mathmale):

nikato's suggested web resource is a good one. http://www.purplemath.com/modules/slope.htm

OpenStudy (anonymous):

@mathmale im lost

OpenStudy (anonymous):

dam. still helping? respect to mathmale

OpenStudy (mathmale):

Sorry, but I was out to dinner with friends. The latest slope is (10-3) / (2-8), or 7/(-6). This slope could also be written -7/6. That's it.

OpenStudy (mathmale):

If you have more questions to discuss on Sunday, please post them.

OpenStudy (anonymous):

@mathmale im hereeeee

OpenStudy (anonymous):

i have 10 more questions to go if you can pleaseeee help me with those i'll pass!!!!

OpenStudy (preetha):

Mathmale, that is an awesome example of help! Many thanks.

OpenStudy (mathmale):

Betty: Can't promise I'll have time to help you with all 10 questions, but I'll help with some. Please look for those 10 questions and pick the ones that appear hardest to you. Choose one of the hardest and post it. Also please respond to that question in some way: by asking questions, by showing preliminary work, or whatever. I want to see what you've done, because that shows me what you know and what you don't know.

OpenStudy (anonymous):

@mathmale im heree i have done some problems but dont know if there right

OpenStudy (nikato):

Just post it. I can try checking them

OpenStudy (anonymous):

i only did two problems so il'l try posting them

OpenStudy (nikato):

Ok

OpenStudy (anonymous):

What is the slope of a line passing through the pair of points (-3/2-5) and (-2 11/2)

OpenStudy (anonymous):

i got 6/-11

OpenStudy (nikato):

Can u rewrite th points? Remember to put commas between the points. And what did u get as ur answers?

OpenStudy (anonymous):

(-3.2,-5) (-2/11,2

OpenStudy (nikato):

So it's 2-(-5) ------ -2/11 -(-3.2)

OpenStudy (anonymous):

its this

OpenStudy (nikato):

R u sure u wrote the problem correctly? It's a really nasty number

OpenStudy (anonymous):

i just sent you the image

OpenStudy (nikato):

Ok, so it's 11/2-(-5) ---------- -2-(-3/2)

OpenStudy (anonymous):

yesss

OpenStudy (nikato):

Can u simplify the top and bottom?